OpenStudy (anonymous):

If dV/dT = 1000e^(0.1t) (a) find v(t) after time t

OpenStudy (anonymous):

I got v(t) by integrating dV/dT v(t) = 10000e^(0.1t) + C

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

unfortunately no lol

OpenStudy (anonymous):

is it dV/dT or dV/dt?

OpenStudy (anonymous):

We need to find C .Plug in t=0 to find C. At time t=0, v(t) = 0 and e^0 = 1 and we get 0 = 10000 +C C = -1000 Your answer is here v= 10000(e^(0.1t) -1)

OpenStudy (anonymous):

i see thanks rita:)

OpenStudy (anonymous):

How do you know that v(0)=0? It's not written in the question.